OpenStudy (anonymous):

In a study of 250 adults, the mean heart rate was 70 beats per minute. Assume the population of heart rates is known to be approximately normal with a standard deviation of 12 beats per minute. What is the 99% confidence interval for the mean beats per minute?

OpenStudy (anonymous):

68.9 – 76.3 70 – 72 61.2 – 72.8 68 – 72

OpenStudy (cj49):

Is the ans 68-72??

OpenStudy (anonymous):

I'm trying to find out for my self

OpenStudy (anonymous):

The confidence interval will look like this: \[\left(\bar{x}-Z_{\alpha/2}\frac{\sigma}{\sqrt n},~\bar{x}+Z_{\alpha/2}\frac{\sigma}{\sqrt n}\right)\] where \[\begin{align*}\bar{x}&:~\text{sample mean}\\ Z_{\alpha/2}&:~\text{critical value for a }(1-\alpha)\times100\%\text{ confidence level}\\ \sigma&:~\text{standard deviation}\\ n&:~\text{sample size} \end{align*}\] Can you identify what's what from the given information?

OpenStudy (anonymous):

not at all

OpenStudy (cj49):

[70 - 2.575.12\div \sqrt{250} ,70+ 2.575.12\div \sqrt{250}\]

OpenStudy (anonymous):

"In a study of 250 adults,...". So \(n=250\). "... the mean heart rate was 70 bpm." So \(\bar{x}=70\). "... with a standard deviation of 12 bpm." So \(\sigma=12\). A 99% confidence level has a critical value of approximately \(Z_{\alpha/2}=2.576\).

OpenStudy (anonymous):

so the answer is b?

OpenStudy (cj49):

B is th ans.